### v1

# Modeling User Intent Beyond Trigger: Incorporating Uncertainty for Trigger-Induced Recommendation

Abstract.
To cater to users’ desire for an immersive browsing experience, numerous e-commerce platforms provide various recommendation scenarios, with a focus on Trigger-Induced Recommendation (TIR) tasks. However, the majority of current TIR methods heavily rely on the trigger item to understand user intent, lacking a higher-level exploration and exploitation of user intent (e.g., popular items and complementary items), which may result in an overly convergent understanding of users’ short-term intent and can be detrimental to users’ long-term purchasing experiences. Moreover, users’ short-term intent shows uncertainty and is affected by various factors such as browsing context and historical behaviors, which poses challenges to user intent modeling. To address these challenges, we propose a novel model called Deep Uncertainty Intent Network (DUIN), comprising three essential modules: i) Explicit Intent Exploit Module extracting explicit user intent using the contrastive learning paradigm; ii) Latent Intent Explore Module exploring latent user intent by leveraging the multi-view relationships between items; iii) Intent Uncertainty Measurement Module offering a distributional estimation and capturing the uncertainty associated with user intent. Experiments on three real-world datasets demonstrate the superior performance of DUIN compared to existing baselines. Notably, DUIN has been deployed across all TIR scenarios in our e-commerce platform, with online A/B testing results conclusively validating its superiority.
Click-Through Rate Prediction; Trigger-Induced Recommendation; Contrastive Learning; User Intent Modeling
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3680065isbn: 979-8-4007-0436-9/24/10ccs: Information systems Recommender systemsccs: Information systems Personalizationccs: Information systems Learning to rank
Figure 1. Recommendation scenarios of an e-commerce platform. Left: Just for You. Middle: Mini Detail. Right: Detail Recommendation.
1. Introduction
Trigger-Induced Recommendation (TIR) (Shen et al., 2022; Xia et al., 2023; Xiao et al., 2024) has recently attracted significant industry attention for its ability to deliver immersive user experiences by modeling the instant history of user clicks on trigger items. For instance, as one of the world’s largest Business-to-Business (B2B) e-commerce platforms, Alibaba.com  [^1] https://www.alibaba.com/ also provides users with many TIR scenarios. As illustrated in Figure 1, by clicking a trigger item in Just for You scenario, users can access some TIR scenarios, e.g., Mini Detail or Detail Recommendation. The primary objective of TIR is to recommend items satisfying users’ needs on conditional of the instant clicking of the trigger item. Nowadays, TIR is playing an increasingly significant role in many industries (Shen et al., 2022; Xie et al., 2021). More than half of active buyers on our website are contributed by TIR among all recommendation scenarios.
Traditional recommendation modeling methods (Cheng et al., 2016; Wang et al., 2018; Li et al., 2020a; Lyu et al., 2020; Pi et al., 2020) often focus on modeling user chronological behaviors with various deep neural networks (Sun et al., 2019; Liu et al., 2019; Xu et al., 2019b), lacking consideration for the trigger item, and thus cannot be directly applied to the TIR scenario. Recently, several trigger-based methods (Shen et al., 2022; Xie et al., 2021; Xiao et al., 2024) have been proposed for the TIR task, which have demonstrated their superiority over traditional methods in the TIR scenario. Nonetheless, user behaviors in real-world e-commerce platforms are more complex (Xu et al., 2020; Qin et al., 2020; Li et al., 2020c; Wang et al., 2021). Figure 2(a) shows an insightful example of users’ subsequent behaviors after clicking a trigger item.
When users clicked on a keyboard, $46.9\%$ of users subsequently purchased the same or similar items (e.g., another keyboard), while $30.8\%$ of them bought trending products (e.g., wireless keyboard) and the rest $22.3\%$ of them bought complementary options (e.g., wireless mouse). This observation reveals that user intent after clicking on a trigger item exhibits a certain degree of uncertainty. This also indicates a need for the platform to offer more precise and comprehensive procurement recommendations.
However, as shown in Figure 2(b), existing TIR methods overly rely on the trigger item, neglecting deeper exploration and exploitation of user intent, which results in an excessively convergent understanding of user intent, and the user will be isolated in a small set of recommended items from the trigger item. For users purchasing computer accessories, products like keyboards and wireless mice are within their intended scope, the overly narrow recommendations can hinder the optimization of users’ long-term purchasing experience. Furthermore, user intent is shaped by a multitude of factors, such as user historical interactions and browsing context. This complexity presents a significant challenge in accurately modeling user intent and capturing the associated uncertainty.
To address these challenges, we propose a novel model in this paper called DUIN  [^2] The source code of DUIN is released at: https://github.com/majx1997/DUIN., short for Deep Uncertainty Intent Network. It primarily comprises three components. To extract explicit user intent based on the trigger, we design an Explicit Intent Exploit Module (EIEM) that uses the contrastive learning paradigm to obtain generalizable and distinguishable intent representations. We also design a Latent Intent Explore Module (LIEM) to effectively explore the users’ underlying intent by leveraging the multi-view relationships between items, applicable to both the trigger item and the target item. An Intent Uncertainty Measurement Module (IUMM) is implemented to model user intent intensity as a Gaussian distribution, capturing the uncertainty associated with user intent. Overall, the contributions of this paper can be summarized as follows:
  • We propose a novel model, Deep Uncertainty Intent Network (DUIN), to address the uncertainty challenges of user intent modeling in Trigger-Induced Recommendation.
  • We introduce an Explicit Intent Exploit Module to extract explicit user intent, and also a Latent Intent Explore Module to explore latent user intent.
  • We design a new Intent Uncertainty Measurement Module to offer distribution estimates and capture the uncertainty of user intent.
  • Extensive experiments on three real-world datasets and an industrial platform demonstrate the superior performance of our model compared to state-of-the-art baselines.
(a) DUIN
(b) Existing TIR methods
Figure 2. Differences in user intent modeling between existing TIR methods and DUIN. The black dashed frame are similar items, red indicates complementary items, and blue denotes trending items. Ratio denotes the proportion of purchasing users. Best viewed in color.
2. Related Work
User behavior Modeling. Existing user behavior modeling approaches (Zhou et al., 2018; Song et al., 2021; Gharibshah et al., 2020; Pi et al., 2019) primarily focus on learning better representation from user behavior sequences and extracting the users’ interests for personalized recommendation. In particular, attention mechanisms(Wang et al., 2019; Tay et al., 2018; Li et al., 2020b; Song et al., 2019; Xu et al., 2019a) have gained popularity in recent years due to their ability to capture complex patterns and temporal dynamics manifested within behavior sequences. For example, DIN (Zhou et al., 2018) utilizes attention mechanisms to activate the interest relation between historical behaviors and the target item. DIEN (Zhou et al., 2019) further integrates a special attention structure with a GRU module to capture the dynamic evolution of user interests. DSIN (Feng et al., 2019) divides user’s sequential behaviors into multiple sessions and employs a self-attention layer to model inner-session interests, alongside a bi-directional LSTM to capture intra-session interests. DMIN (Xiao et al., 2020) explores user’s multiple diverse interests and introduces a multi-head self-attention layer and a multi-interest extractor layer to represent user’s interests by multiple vectors. These attention-based methods have demonstrated feasible performance across various scenarios. Although the aforementioned recommendation approaches can be directly applied to TIR scenarios, their inability to account for modeling user’s immediate intent representing by the trigger item leads to challenges in accurately estimating user intent in real-world industrial scenarios.
Trigger-Induced Recommendation. Though above traditional solutions can be rigidly applied to TIR scenarios, the lack of considering the trigger item usually make them get suboptiomal results in these scenarios. Recently, various methods have been proposed to target for the TIR problem. DIHN (Shen et al., 2022) is one pioneer in introducing the TIR task, and utilizes a deep network to predict the user’s real intent regarding the trigger item. DIAN (Xia et al., 2023) proposes an intention-aware network to extract the user’s intention and balance the outcomes of trigger-free and trigger-based recommendations. DEI2N (Xiao et al., 2024) introduces a user instant interest modeling layer to forecast the dynamic change in the intensity of instant interest when the user clicks on a trigger item and scrolls down. These existing works have showcased their superiority over traditional methods in the TIR scenario. However, the heavy reliance of these recently proposed methods on trigger items limits their capacity to capture the intricate and higher-level relationships in user behavior. As a result, they may fail to fully reveal the extent of user’s true intent, potentially leading to an overly convergent understanding of user intent. This limitation could adversely affect users’ long-term experiences.
3. Method
Figure 3. The architecture of DUIN consists of three modules, EIEM, LIEM and IUMM.
The overall architecture of the proposed DUIN is illustrated in Figure 3, which mainly consists of three modules: Explicit Intent Exploit Module (EIEM), Latent Intent Explore Module (LIEM), and Intent Uncertainty Measurement Module (IUMM). Before diving into the details of these modules, we first formulate the TIR task and introduce some necessary notations.
3.1. Problem Setup
Let $\mathcal{U}=\{u_{1},u_{2},...,u_{M}\}$ and =$\mathcal{I}=\{i_{1},i_{2},...,i_{N}\}$ represent the user set and the item set, respectively. Each user has a behavior sequence $\mathcal{S}^{u}=\{i_{1}^{u},i_{2}^{u},...,i_{T}^{u}\}$, where $M$, $N$, and $T$ denote the number of users, items, and the length of behavior sequences, respectively. We denote user profile features as $u^{p}$ and context features as $u^{c}$. Based on these notations, given the trigger item $i^{tr}$, the primary objective of TIR is to predict the probability $\hat{y}$ of user $u$ potentially interacting with the target item $i^{ta}$ as follows:
(1) $\hat{y}=\mathcal{F}_{\theta}(\mathbf{E}(u^{p}),\mathbf{E}(u^{c}),\mathbf{E}(i^{tr}),\mathbf{E}(i^{ta}),\mathbf{E}(\mathcal{S}^{u})),$
where $\mathcal{F}_{\theta}$ represents the prediction function we aim to learn. The notation $\mathbf{E}(\cdot)\in\mathbb{R}^{d}$ denotes the embedding layer (Cheng et al., 2016), which transforms raw features into dense vectors with dimension $d$.
3.2. Deep Uncertainty Intent Network
3.2.1. Explicit Intent Exploit Module
User chronological behaviors exhibit causality and convey explicit intent information, including aspects users are inclined to interact with, such as brand, theme, and category. The trigger item merely represents a partial expression of user explicit intent. In DUIN, we introduce the Explicit Intent Exploit Module to learn user explicit intent.
Typically, items with identical attributes exhibit a high degree of similarity. For instance, in the e-commerce scenario, products that belong to the same category often share similar intrinsic quality or interacted uses. In user behavior sequences, we consider items that share attributes with the trigger item express the user explicit intent. Given a trigger item $i^{tr}$, we retrieve items from the behavior sequence $\mathcal{S}^{u}$ that have the same attribute as $i^{tr}$. Together with the trigger item, these form an explicit intent sequence $\mathcal{S}^{e}$:
(2) $\mathcal{S}^{e}=\{i^{tr},i_{1}^{e},...i_{l}^{e},...,i_{L}^{e}\},$
where $i_{l}^{e}$ denotes items that share the same attribute with the trigger item $i^{tr}$. $L$ represents the number of $i_{l}^{e}$.
Various attention-based methods (Zhou et al., 2018; Kang and McAuley, 2018) have significantly facilitated user behavior modeling. However, recent research (Qiu et al., 2022; Xie et al., 2022; Chen et al., 2022) indicates that user representations generated by these methods tend to degenerate into an anisotropic shape, which may result in high semantic similarities among representations. To address this issue, inspired by recent advances of contrastive learning paradigm in the computer vision (Chen et al., 2021, 2020), based on the explicit intent sequence, we apply contrastive learning to obtain discriminative and high-quality user explicit intent representations. Formally, the explicit intent contrastive learning task can be defined as:
(3) $f_{\theta}=\operatorname*{arg\,min}\limits_{f_{\theta}}\mathcal{L}_{ssl}(f_{\theta}(\mathcal{S}^{e}),f_{\theta}(\widetilde{\mathcal{S}}^{e}_{+})),$
where $\mathcal{L}_{ssl}$ is the contrastive loss function. $f_{\theta}$ is the shared encoder. $\widetilde{\mathcal{S}}^{e}_{+}$ is the augmented view of the explicit intent sequence $\mathcal{S}^{e}$.
In practice, we mask the explicit intent sequence $\mathcal{S}^{e}$ with a certain probability to get the augmented view:
(4) $\widetilde{\mathcal{S}}^{e}_{+}\sim\mathcal{A}(\mathcal{S}^{e},\gamma),$
where $\mathcal{A}(\cdot)$ denotes the augmentation operator and $\gamma$ is the mask probability. The idea behind this is that user explicit intent is relatively stable during a period of time. Therefore, though some of the items are masked, the explicit intent information is still retained and should be considered as the positive view. Thus, we treat $(\mathcal{S}^{e},\widetilde{\mathcal{S}}^{e}_{+})$ as positive view pairs and other augmented examples within the same batch as negative views $\widetilde{\mathcal{S}}^{e}_{-}$.
Furthermore, we adopt the shared encoder $f_{\theta}$ to extract the user explicit intent representation $\mathcal{H}_{ei}=f_{\theta}(S^{e})$. We optimize $f_{\theta}$ via a contrastive loss function (Chen et al., 2020), which can be formulated as:
(5) $\mathcal{L}_{ssl}=-log\frac{e^{sim(\mathcal{H}_{ei},\mathcal{H}_{ei}^{+})/\tau}}{e^{sim(\mathcal{H}_{ei},\mathcal{H}_{ei}^{+})/\tau}+\sum_{i=1}^{2B-1}e^{sim(\mathcal{H}_{ei},\mathcal{H}_{ei}^{-})/\tau}},$
where $\mathcal{H}_{ei}^{+}$ and $\mathcal{H}_{ei}^{-}$ indicate the augmented positive view and the negative views of $\mathcal{H}_{ei}$ respectively. $sim(\cdot)$ denotes the cosine similarity function. $B$ is the batch size. $\tau$ is the temperature parameter.
It is worth mentioning that EIEM also introduces a feature interaction module to learn the explicit interaction relationship between the target item $i^{ta}$ and the user explicit intent $\mathcal{H}_{ei}$, the output representation $\mathcal{H}_{i}$ is calculated as:
(6) $\mathcal{H}_{i}=MLP(\mathcal{H}_{ei},\mathbf{E}(i^{ta}),\mathcal{H}_{ei}\odot\mathbf{E}(i^{ta}),\mathcal{H}_{ei}-\mathbf{E}(i^{ta})),$
where $\odot$ denotes the Hadamard product.
3.2.2. Latent Intent Explore Module
As aforementioned, users have various latent intents after clicking on the trigger item, such as seeking similar items, popular items, or complementary items. Previous TIR methods (Shen et al., 2022; Xia et al., 2023; Xiao et al., 2024) lack exploration of user latent intent and overlook the potential information related to user intent within user behavior sequences. In DUIN, we propose the Latent Intent Explore Module to reveal user latent intent.
We consider users with similar intents usually exhibit similar behaviors, thus, we aggregate the temporal directed behaviors of multiple users to form a fundamental probabilistic graph structure $\mathcal{G}$. The nodes $\mathcal{V}$ in the graph contain items and their attributes, and we will connect two sequentially occurring nodes to refine collaborative relationships (Wu et al., 2019; Chang et al., 2021; Min et al., 2022).
Having obtained the finer-grained collaborative patterns in graph $\mathcal{G}$, suppose there is an item $i$ with attribute $\phi$ and an item $j$ with attribute $\psi$, we construct multi-view relationships between item $i$ and item $j$ as follows:
(7) $\mathcal{R}_{ij}^{t}$ $=\mathcal{C}_{i\rightarrow j}(\{(\mathcal{V}_{i},\mathcal{V}_{j})|(\mathcal{V}_{i},\mathcal{V}_{j})\in\mathcal{G}\}),$
(8) $\mathcal{R}_{ij}^{c}$ $=\mathcal{C}_{\phi\rightarrow\psi}(\{(\mathcal{V}_{\phi},\mathcal{V}_{\psi})|(\mathcal{V}_{\phi},\mathcal{V}_{\psi})\in\mathcal{G}\}),$
(9) $\mathcal{R}_{ij}^{p}$ $=\mathcal{C}_{\phi\rightarrow j}(\{(\mathcal{V}_{\phi},\mathcal{V}_{j})|(\mathcal{V}_{\phi},\mathcal{V}_{j})\in\mathcal{G}\}),$
where $\mathcal{V_{*}}$ denotes the note in graph $\mathcal{G}$. $\mathcal{C_{*}}$ is the aggregation operator that records two nodes’ co-occurrence frequency.
The high-order connections between items can be implicitly captured by the aforementioned multi-view relationships. Specifically, we use $\mathcal{C}_{i\rightarrow j}$ to calculate the co-occurrence frequency between item $i$ and item $j$ and it is straightforward to use $\mathcal{R}_{ij}^{t}$ to represent their directed transition relationship. Similarly, the collaborative relationship of attributes, denoted by $\mathcal{R}_{ij}^{c}$, represents the complementary relationship between item $i$ and item $j$. $\mathcal{R}_{ij}^{p}$ is utilized to characterize the popularity of item $j$ under attribute $\phi$.
For item $i$ and item $j$, to better encode the information from multi-view relationships between them, we define the latent intent relevance score $\Pi_{ij}$ to quantify their potential connections:
(10) $\Pi_{ij}=Sigmoid(MLP(\mathbf{E}(\mathcal{R}_{ij}^{t}),\mathbf{E}(\mathcal{R}_{ij}^{p}),\mathbf{E}(\mathcal{R}_{ij}^{c}))).$
Motivated by previous work (Shen et al., 2022; Xiao et al., 2024), we employ Multi-Head Self-Attention (MHSA) (Vaswani et al., 2017) to obtain the refined behavior representation $\mathbf{E}(\mathcal{S}^{u})$. Next, we apply two MHSA modules to extract user latent intent with respect to the target item and the trigger item respectively. More importantly, we propose a simple yet effective method to enable that the attention mechanism aware of the prior multi-view relationships among items. Specifically, for each behavior item $i_{*}^{u}\in\mathcal{S}^{u}$, we separately calculate the latent intent relevance score with both the trigger item and the target item according to equation 10, and we also modify the attention calculation scheme as follows:
(11) $\mathcal{H}_{li}^{tr}$ $=Attention(\mathbf{E}(i^{tr}),\Pi^{tr}\mathbf{E}(\mathcal{S}^{u}),\Pi^{tr}\mathbf{E}(\mathcal{S}^{u})),$
(12) $\mathcal{H}_{li}^{ta}$ $=Attention(\mathbf{E}(i^{ta}),\Pi^{ta}\mathbf{E}(\mathcal{S}^{u}),\Pi^{ta}\mathbf{E}(\mathcal{S}^{u})),$
where $\mathcal{H}_{li}^{tr}$ and $\mathcal{H}_{li}^{ta}$ are user latent intent representation with respect to the trigger item and the target item respectively. $\Pi^{tr}$ represents the latent intent relevance score between each behavior item and the trigger item. $\Pi^{ta}$ denotes the relevance score between each behavior item and the target item.
3.2.3. Intent Uncertainty Measurement Module
In the TIR task, supposing a continuous mapping space $\mathcal{X}\rightarrow\mathcal{Z}$, where $x_{u}\in\mathcal{X}$ denotes user personalized information, including the trigger item $i^{tr}$, context features $u^{c}$, and profile features $u^{p}$. $z_{u}\in\mathcal{Z}$ represents user intent intensity. As user intent evolves and browsing context changes, the intensity of user intent after clicking on the trigger item becomes uncertain to some degree. Therefore, in DUIN, we consider the user interaction behavior as an uncertain event influenced by the user intent intensity. We employ the Intent Uncertainty Measurement Model to model user intent intensity as a distribution rather than a static value (Shen et al., 2022; Xiao et al., 2024), which can cover a broader space of user intent and infuse the recommendation results with novelty and uncertainty.
Specifically, we define the user intent intensity $z_{u}$ of each user as a Gaussian distribution $\mathcal{N}$, which is mathematically defined as:
(13) $\mathit{p}(z_{u}|x_{u})=\mathcal{N}(z_{u};\mu_{u},\Sigma_{u}\textbf{{I}}),$
where $\mu_{u}$ and $\Sigma_{u}$ denote the mean and the variance of the Gaussian distribution, respectively. $\mu_{u}$ represents the predicted user intent intensity and $\Sigma_{u}$ can be regarded as the uncertainty of $\mu_{u}$. Here we only consider a diagonal matrix for simplicity. In practice, we employ two independent networks to obtain $\mu_{u}$ and $\Sigma_{u}$:
(14) $\mu_{u}$ $=f_{\theta}^{\mu}(x_{u}),$
(15) $\Sigma_{u}$ $=softplus(f_{\theta}^{\Sigma}(x_{u})),$
where $f_{\theta}^{\mu}$ and $f_{\theta}^{\Sigma}$ are implemented as two fully-connected layers. $softplus$ is an activation function (Glorot et al., 2011) and constrains the output to always be positive.
Finally, we consider $z_{u}$ as a measure of the user’s intent intensity towards the trigger item. We use it to integrate the user latent intent features outputted by LIEM, which can be formulated as:
(16) $\mathcal{H}_{li}=(z_{u}\odot\mathcal{H}_{li}^{tr};(1-z_{u})\odot\mathcal{H}_{li}^{ta}).$
3.3. Prediction and Optimization
Having obtained the user intent representations $\mathcal{H}_{ei}$, $\mathcal{H}_{i}$ and $\mathcal{H}_{li}$ from both EIEM and LIEM, similar to equation 1, we concatenate these intent representations with other features, including context features $u^{c}$, profile features $u^{p}$, the target item $i^{tr}$, and others. Subsequently, the concatenated features are fed into MLP (Zhou et al., 2018) to estimate the likelihood $\hat{y_{i}}\in[0,1]$ of user interaction. We adopt the binary cross entropy loss (Zhou et al., 2018; Shen et al., 2022; Wang et al., 2017) for model optimization:
(17) $\mathcal{L}_{ctr}=-\frac{1}{N}\sum\limits_{i=1}^{N}(y_{i}log(\hat{y_{i}})+(1-y_{i})log(1-\hat{y_{i}})),$
where $N$ denotes total size of the training set, $y_{i}$ is the ground-truth. We also adopt a multi-task learning strategy to jointly optimize the original prediction task and the contrastive learning objectives. The final optimization function is defined as:
(18) $\mathcal{L}_{final}=\mathcal{L}_{ctr}+\alpha\cdot\mathcal{L}_{ssl},$
where $\alpha$ is a hyperparameter to balance weight of contrastive loss.
4. Evaluation
4.1. Experimental Setup
4.1.1. Datasets.
Three real-world datasets are used in experiments. Detailed statistics are summarized in Table 1.
  • Alibaba.com  [^3] https://www.alibaba.com/. As there is no public dataset released for the TIR task. We collect interaction behavior logs from the TIR scenarios on our e-commerce application to serve as industrial production dataset, named Alibaba.com. The dataset contains about 370 thousand users and 5.2 million records of browsing and clicking behavior data over 3 days. We select data from the first 80% in chronological order as the training set, the next 10% as validation set, while the rest 10% as test set.
  • Alimama  [^4] https://tianchi.aliyun.com/dataset/dataDetail?dataId=56. It is a representative e-commerce dataset provided by Alimama advertising platform. Since it lacks logs of trigger items, following the data processing strategy by  (Shen et al., 2022), for each sample, we take the latest item clicked by user within four hours as the trigger item. Finally, the dataset contains 8.55 million samples from half a million users.
  • ContentWise (Pérez Maurera et al., 2020). Different from Alibaba.com and Alimama datasets, the dataset is a popular multimedia recommendation dataset, which is collected from an Over-The-Top Media service and includes 2.5 million behavior data from approximately 26 thousand users. Similar to the Alimama dataset, we consider the latest clicked item by a user within eight hours as the trigger item.
Table 1. Datasets statistics.
Dataset #Users #Items #Attributes #Samples
Alibaba.com 373,852 4,715,150 6,736 5,200,000
Alimama 500,000 846,812 12,978 8,552,702
ContentWise 26,186 1,268,988 117,693 2,585,070
4.1.2. Compared Methods.
We include three groups of baseline methods for comparison. We firstly compare DUIN with some popular user behavior modeling methods: WDL (Cheng et al., 2016), DIN (Zhou et al., 2018), DIEN (Zhou et al., 2019), DMIN (Xiao et al., 2020). Furthermore, following (Shen et al., 2022; Xia et al., 2023; Xiao et al., 2024), we upgrade previous methods by incorporating the trigger item, forming the second group for a fair comparison:
  • WDLTIR concatenates the trigger item with other features and feds them into the deep side of the network.
  • DINTIR employs an attention mechanism to activate related user behaviors with respect to the trigger item.
  • DIENTIR utilizes GRU with an attentional gate to extract the evolved interests that are related to the trigger item.
  • DMINTIR applies interest extraction layers to obtain user multiple interests that are related to the trigger item.
The third group involves several state-of-the-art methods focusing on the TIR task: DIHN (Shen et al., 2022), DIAN (Xia et al., 2023), DEI2N (Xiao et al., 2024).
4.1.3. Implementation details.
We employ the Adam optimizer for all models with a learning rate of 0.001 and a batch size of 256. Each model is trained from scratch without any pre-training. The maximum lengths of the user sequences in the Alibaba.com, Alimama, and ContentWise datasets are 20, 20, and 30, respectively. We set the maximum length of user explicit intent sequences to 10 for the Alibaba and Alimama datasets and to 20 for the ContentWise dataset. We select category and brand as item attributes for e-commerce dataset and multimedia dataset respectively. We construct probability graph by linking each item to its four subsequent items. We utilize the Transformer (Vaswani et al., 2017) architecture as the encoder $f_{\theta}$, the number of attention heads is set to 8 for all. The MLP within the EIEM and IUMM are configured with hidden layer sizes of 144 and 72. The dimensions of the final hidden layers are set at 200 and 80, respectively. For contrastive learning, the hyperparameters $\tau$, $\gamma$, and $\alpha$ are assigned values of 0.1, 0.5, and 1, respectively.
4.1.4. Evaluation Metrics.
In offline experiments, we use commonly used metrics AUC and RelaImpr (Zhou et al., 2019; Xiao et al., 2020; Feng et al., 2019), which are widely recognized evaluation metrics in CTR tasks. We conduct each experiment five times and report the average results. All other experimental and parameter settings remain consistent with  (Shen et al., 2022; Xiao et al., 2024) to ensure fair comparison. One-sided Wilcoxon rank-sum p-value is calculated to quantify the significance of the improvement achieved by DUIN (Gehan, 1965). In online A/B testing, we use CTR and Conversion Rate (CVR) as online metrics to assess the efficacy of DUIN.
Table 2. Performance comparison of our method with competitors on three real-world datasets. Bold values indicate the best result in each column, while underlined values indicate the second best result.
Model Alibaba.com Alimama ContentWise
AUC RelaImpr AUC RelaImpr AUC RelaImpr
WDL $0.6096\pm 0.0019$ $-0.99\%$ $0.6062\pm 0.0008$ $-7.97\%$ $0.9469\pm 0.0003$ $-7.28\%$
DIN $0.6042\pm 0.0016$ $-5.87\%$ $0.6154\pm 0.0007$ $0.00\%$ $0.9774\pm 0.0002$ $-0.95\%$
DIEN $0.6047\pm 0.0025$ $-5.42\%$ $0.6155\pm 0.0005$ $0.09\%$ $0.9779\pm 0.0013$ $-0.85\%$
DMIN $0.6107\pm 0.0011$ $0.00\%$ $0.6154\pm 0.0002$ $0.00\%$ $0.9820\pm 0.0002$ $0.00\%$
WDLTIR $0.7412\pm 0.0014$ $111.89\%$ $0.6075\pm 0.0018$ $-6.84\%$ $0.9713\pm 0.0004$ $-2.22\%$
DINTIR $0.7425\pm 0.0021$ $119.06\%$ $0.6155\pm 0.0015$ $0.09\%$ $0.9803\pm 0.0019$ $-0.35\%$
DIENTIR $0.7419\pm 0.0019$ $118.52\%$ $0.6157\pm 0.0004$ $0.26\%$ $0.9796\pm 0.0015$ $-0.50\%$
DMINTIR $0.7454\pm 0.0007$ $121.68\%$ $0.6157\pm 0.0003$ $0.26\%$ $0.9822\pm 0.0003$ $0.04\%$
DIHN $0.7462\pm 0.0006$ $122.40\%$ $0.6166\pm 0.0008$ $1.04\%$ $0.9786\pm 0.0012$ $-0.75\%$
DIAN $0.7480\pm 0.0016$ $124.03\%$ $0.6168\pm 0.0002$ $1.21\%$ $0.9764\pm 0.0003$ $-1.16\%$
DEI2N $0.7671\pm 0.0012$ $141.28\%$ $0.6180\pm 0.0005$ $2.25\%$ $0.9840\pm 0.0002$ ${0.41\%}$
DUIN $\textbf{0.7782}\pm\textbf{0.0014}$* 151.31% $\textbf{0.6194}\pm\textbf{0.0003}$* 3.47% $\textbf{0.9881}\pm\textbf{0.0004}$* 1.27%
  • $\ast$
    Asterisks represent where DUIN’s improvement over compared methods is significant (one-sided rank-num p-value ¡0.01).
4.2. Offline Experiments
4.2.1. Overall Comparison
The overall performance of all methods across three datasets is summarized in Table 2. From the results, we can observe that traditional CTR methods deliver the poorest results compared with other methods, indicating their inadequacy for the TIR task. By incorporating the trigger item, upgraded versions of these traditional methods demonstrate much improved performance. For example, DMINTIR has a AUC gains $22.1\%$ over its original version for Alibaba.com dataset. It shows that an elaborative modeling of the trigger item is necessary. Meanwhile, existing TIR methods, DIHAN, DIAN and DEI2N, employ carefully designed networks to further surpass the performance of previous upgraded traditional approaches. Notably, our proposed DUIN achieves the best performance across all three datasets. The results show that the AUC gains of our method DUIN over the existing strongest competitor DEI2N are $1.45\%$, $0.22\%$ and $0.41\%$ for Alibaba.com, Alimama and ContentWise datasets respectively. It validates the efficacy of considering explicit and latent intents and the corresponding uncertainties in TIR scenarios.
Table 3. Ablation experimental results on Alibaba.com.
Model Module Alibaba.com
EIEM LIEM IUMM SSL SII AUC RelaImpr
1 $\textbf{0.7782}\pm\textbf{0.0014}$ $\textbf{0.0}\%$
2 $0.7637\pm 0.0011$ $-5.21\%$
3 $0.7653\pm 0.0013$ $-4.63\%$
4 $0.7536\pm 0.0009$ $-8.84\%$
5 $0.7730\pm 0.0015$ $-1.87\%$
6 $0.7661\pm 0.0007$ $-4.35\%$
4.2.2. Ablation Study.
As Alibaba.com is collected from real-world TIR scenarios, we conduct several ablation experiments on it to further analyze the effectiveness of each module. Table 3 shows the experimental results, where SSL represents the contrastive learning framework in EIEM, and SII denotes that DUIN utilizes a static intent intensity to replace IUMM by following (Shen et al., 2022; Xiao et al., 2024). Compared to Model 2-4, Model 1 achieves the highest AUC results, indicating that all modules are effective for DUIN. The superior performance of Model 1 over Model 4 shows that modeling the user intent intensity is critical in the TIR task. The ablation results of Models 1 and 5 confirm that contrastive learning can help improve the performance of DUIN. By comparing Models 1 and 6, we can conclude that modeling user intent intensity as a distribution, rather than a static value is a more effective representation method.
4.2.3. Hyper-parameters Analysis.
We further qualitatively analyze the impacts of different hyper-parameters in DUIN, which are shown in Figure 4(a). Overall, DUIN demonstrates relatively robust performance across parameter variations. We also observe that the performance of DUIN deteriorates when $\gamma$ is too large, probably because the excessive loss of interactive behaviors affects the modeling of user explicit intent.
(a) Hyper-parameters Analysis
(b) Time Analysis
Figure 4. Parameters and Time Analysis on Alibaba.com.
4.2.4. Time Analysis.
We also conduct an extra experiment to test the efficiency of DUIN. To make comparisons fairly, we conduct experiments on machines with identical hardware configurations, each of which has 1 NVIDIA Tesla V100 GPU with Intel(R) Xeon(R) Platinum 8163 CPU @ 2.50GHz and 330 GB memory. We report the time consumption for one epoch of training and testing for each model, respectively. As shown in Figure 4(b), compared to the existing TIR methods, DUIN slightly increases the training time. However, since contrastive learning is only utilized in the training stage, DUIN does not introduce additional prediction time cost.
4.3. Online A/B Testing
4.3.1. Online Serving Architecture
Figure 5 depicts the online serving architecture of DUIN on our e-commerce platform. The architecture is primarily divided into two main stages: i) The offline stage focuses on training the DUIN model. In practice, we collect 60 days of user interaction logs from our platform and process these logs into training samples using the big data platform MAXCOMPUTE. Once the training samples are prepared, we utilize the Algorithm One Platform (AOP) to train our proposed DUIN model. After training is complete, the model is deployed for online serving. ii) During the online stage, The Personalization Platform (TPP) handles user requests by parsing them to obtain the real-time clicked trigger item and context features of the current browsing session. TPP then requests All Basic Feature Service (ABFS) to retrieve user profile features and historical behaviors. The Basic Engine (BE) recalls the top-M most relevant candidate items for the user, where M is typically less than two thousand. Our DUIN model is deployed on the Real-Time Prediction (RTP) platform to score these top-M candidate items. The top-K items with the highest scores are then shown to the user. Usually, the $K$ is set to 10-20. In addition, RTP supports deploying multiple models, ensuring fairness in online A/B testing.
4.3.2. Online A/B Testing Results
As shown in Table 4, we conduct online A/B testing to evaluate DUIN in Detail Recommendation and Mini Detail, denoted as TIR1 and TIR2 respectively. The Daily Active Users (DAU) for the two TIR scenarios are approximately 700K and 300K, respectively. We choose DEI2N (Xiao et al., 2024) as the baseline, as it performs the second-best in offline tests. We conduct 14-day experiments in the two scenarios, and the experimental results are statistically significant according to an unpaired t-test. Results show that DUIN consistently outperforms DEI2N in terms of CTR and CVR. Furthermore, since the exposed items can be classified into two categories—those that belong to the same category as the trigger item and those that belong to a different category, CTR can also be segmented into Same-category CTR (S-CTR) and Cross-category CTR (C-CTR). DUIN also shows significant improvements under both S-CTR and C-CTR metrics, indicating its superior intent representation capabilities. From the perspective of cross-category exposure ratio (CCER), DUIN alleviates the problem of over-convergence in intention understanding and provide users with more diverse and novel recommendations. Due to DUIN’s excellent online performance, we have deployed it in major TIR scenarios on the Alibaba.com platform.
Table 4. Online A/B testing results.
Scenario Model CTR CVR S-CTR C-CTR CCER
TIR 1 DEI2N 4.61% 3.96% 5.01% 2.70% 41.09%
DUIN 4.75% 4.08% 5.25% 2.81% 46.81%
TIR 2 DEI2N 2.73% 1.59% 2.81% 2.92% 27.75%
DUIN 2.82% 1.64% 2.88% 3.01% 28.50%
4.3.3. Case study
We illustrate the effectiveness of DUIN through two real cases on the Alibaba.com platform. As depicted in Figure 6, the upper section illustrates that when a user, for instance, a headphone retailer, engages in our TIR scenario. When he/she clicks on the Blue Bluetooth Headphones, DUIN suggests items that are akin to the trigger item. At the same time, DUIN suggests trending items and complementary options to the user, like the Wired Earphone and Headset. This shows that DUIN not only recommends items similar to the trigger item but also accurately predicts the user’s intended scope, offering more precise and comprehensive procurement recommendations. The lower section of Figure 6 shows another real-world scenario. After a clothing wholesaler clicked on the Ladies T-Shirt, DUIN not only suggests some Ladies T-Shirts similar to the trigger item but also recommends Mini Skirts that can be easily coordinated with the Ladies T-Shirt, as well as Cotton Tank Tops that are currently trending on the platform.
Figure 5. The online serving architecture of DUIN.
Figure 6. Two real-world cases. Black dashed boxes are similar items, red dashed boxes indicates complementary items, and blue dashed boxes indicates trending items.
5. Conclusion
In this paper, we introduce a novel TIR model, called Deep Uncertainty Intent Network (DUIN), to model both explicit and latent user intent with uncertainty based on user-triggered items. The proposed DUIN consists of three modules: i) Explicit Intent Exploit Module, which aims to obtain generalizable and distinguishable intent representations; ii) Latent Intent Explore Module, designed to explore the user’s underlying intent; and iii) Intent Uncertainty Measurement Module, tasked with modeling user intent intensity as a distribution to capture the uncertainty associated with user intent. Extensive experiments demonstrate that DUIN surpasses various representative methods, achieving state-of-the-art performance. Moreover, DUIN has been deployed across all TIR scenarios within our Alibaba.com platform, and online A/B testing has conclusively verified the superiority of our approach.